The sum of a certain even no.and the fourth even no. After it is 48 find the number.
Answer by ankor@dixie-net.com(22740)

The sum of a certain even no. and the fourth even no. after it is 48.
find the number.
n + (n+8) = 48
2n = 48 - 8
2n = 40
n = 40/2
n = 20 is the number
:
:
Check
20 + 28 = 48